OpenStudy (anonymous):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9)

OpenStudy (anonymous):

The slope between two points is equal to the difference in their y values divided by the difference in their x values: \[\frac{ \Delta y }{ \Delta x } = \frac{ 9--3 }{ 2-2 } = \frac{ 12 }{ 0 } = Undefined\]